OpenStudy (anonymous):

Simplify the following expression. square root of -121-24

OpenStudy (anonymous):

\[\sqrt{-121-24}\] \[\sqrt{-145}\] \[\sqrt{-1}\sqrt{145}\] \[i\sqrt{145}\]

OpenStudy (anonymous):

can you figureo ut what youcan pull out of the square root now?

OpenStudy (anonymous):

how do you get 145 it said that you times your radical at the end to get the final answer

OpenStudy (anonymous):

the square root of -121 is 11 but i sqaure root of -24 is 4*6 can you help me please

OpenStudy (anonymous):

so your question is \[\sqrt{-121}-\sqrt{-24}\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i think i finally got it is the answer 2√6 -24

jhonyy9 (jhonyy9):

so sqrt(-121) =

jhonyy9 (jhonyy9):

and sqrt(-24) =

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you have to simplify

jhonyy9 (jhonyy9):

sqrt(-121) = sqrt((- (11*11)) =

jhonyy9 (jhonyy9):

= 11 sqrt(-1) =

jhonyy9 (jhonyy9):

so you know that sqrt(-1)=i so than sqrt(-121) = 11i ok ?

jhonyy9 (jhonyy9):

so can you continue it with sqrt(-24) ?

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

not is right there is a minus too

OpenStudy (anonymous):

what is not right

jhonyy9 (jhonyy9):

so there is sqrt(-121) - sqrt(-24) = yes ,this is the exercise and little up was evalueted sqrt24 = 2sqrt6 ok ?